Contract the expressions. That is, use the properties of logarithms to write each expression as a single logarithm with a coefficient of 1.
ln (3) - 2 ln (4) + ln (8)
ln (3) - 2 ln (4+8)
ln *3) - 2 (ln (4) +ln (8)

Answer by stanbon(57347)

write each expression as a single logarithm with a coefficient of 1.
ln (3) - 2 ln (4) + ln (8) = ln[3*8/4^2] = ln[24/16] = ln(3/2)
----------------------------------------------
ln (3) - 2 ln (4+8) = ln[3/12^2] = ln[3/144]
--------------------------------
ln (3) - 2 ln (4) +ln (8) = ln[3*8/4^2] = ln[24/16] =========================================================
Cheers,
Stan H.